OpenStudy (anonymous):

1. Fill in the blanks: a. The graph of a quadratic function is a symmetric, U-shaped curve called a __________________. b. The highest or lowest point of the graph of a quadratic function is called the _________________. c. Standard Form of a quadratic equation is: ____________________________. d. The graph is symmetric about the ___________________ which is x= e. Vertex form of a quadratic equation is ___________________and the vertex is at the point ( , ). heres some of my answers = a/ Parabola b/ vertex c/ y = ax2 + bx + c d/ ? e/ ?

OpenStudy (anonymous):

Parabola Vertex y=ax^2+bx+c Vertex or A.O.S. (Axis of symmetry) occurs at Max/min of the parabola The vertex form of a quadratic function is given by f (x) = a(x - h)2 + k, where (h, k) is the vertex of the parabola

OpenStudy (anonymous):

Hope this helps

OpenStudy (anonymous):

Yes thanks so much !!

OpenStudy (anonymous):

can you help with a few more

OpenStudy (anonymous):

The length of a rectangular yard is 5 feet longer than its width, x. The area of the yard is 104 sq ft. Use a quadratic to solve for the width. Use completing the square to convert from standard form into vertex form: f(x)=x^2=8x=10